OpenStudy (anonymous):

how many positive odd integers less than 70,000 can be represented using the digits 2,3,4,5,and 6

OpenStudy (anonymous):

greaaaaat so nobody knows:(

OpenStudy (anonymous):

just looked at it now, thinking

OpenStudy (anonymous):

you mean as a arithmetic combination of these numbers or does it mean like you can use only 2,3,4,5,6 to represent the number? like 23 is a valid number but 37 is not because it uses 7?

OpenStudy (anonymous):

I'm sure there's some fancy statistics way to do this, but thinking about it the long way... For one-digit numbers: only 3 works so 1 For two digit numbers: 3 still has to be the second digit, but any of the 5 numbers can be first, so 5 For three digit numbers: 25, by similar logic for four: 125 for five: 625 for six: 3125 so 1 + 5 + 5^2 + 5^3 + 5^4 + 5^5 = 3906 I'm not totally sure of my method. That make good sense?

OpenStudy (anonymous):

Oh woops, went one exponent too far. Six-digit numbers don't qualify. so 1 + 5 + 5^2 + 5^3 + 5^4 = 781

OpenStudy (anonymous):

but 5 is an odd number too :)

OpenStudy (anonymous):

Right...somehow didn't even see the 5. Apologies. What would that do?...it would mean that both 3 and 5 could be the last digit, and make for twice as many possibilities, right?

OpenStudy (anonymous):

not sure, but I think not

OpenStudy (anonymous):

I give up, it is too late here :) tomorrow I will give it another go if nobody solved it till then

OpenStudy (anonymous):

I'm relatively confident that it's 781*2 = 1562, but I know my work is pretty sloppy.